Davis Ltd.’s total assets and total liabilities were $27,000 and $18,000 respectively at the beginning of the current year, and $31,000 and $20,000 respectively at the end of the current year. If the company issued no common shares during the year and earned $4,000 in net income, how much did it declare in dividends?
Answer #1

Stockholder's equity = Total assets-Total liabilities

Beginning = 27000-18000 = 9000

Ending = 31000-20000 = 11000

Dividend = Beginning equity+Net income-Ending equity

             = 9000+4000-11000

Dividend = 2000